Home‎ > ‎Crew 165‎ > ‎

0310-EVAReport

EVA Report:

 EVA #3

 Crew members: Executive Officer Sullivan, Biologist IIzuka, GreenHab Officer Hiraki and Journalist Mori

 Location: Around the Hab (12S-518155E-425052N), Sentinel Hill (12S-519420E-4250245N)

 Time: Departure 10:55 a.m., Return 13:55 p.m.

 Duration: 3 hours

 Purposes:

 NorCal Rover testing (retrying with additional batteries because batteries
died so quickly in EVA 2), biological sampling, photo/video

 Summary:

 *(11:15) GreenHab resupply shipment unloaded from docking
 station and relocated inside for assembly.

 *(11:30) Outside static tank was resupplied with water from our
 Galactic Water Provider.

 *(11:50) Rover was returned to Cycled Engineering Airlock for
 diagnostics after wireless connection was lost.


 *(12:00) ATV 4 parking brake repaired, five fully functioning
 ATV’s waiting refueling supplies.

 *(12:30) Collect soil sample and photo and video at Sentinel
 Hill.

 *(13:40) Collect soil sample at septic area.